# Variations in Relevance Judgments and the Shelf Life of Test Collections

### v1

@@ Abstract.
ccs: Information systems Evaluation of retrieval resultscopyright: acmcopyrightjournalyear: 2025conference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2024; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/XX/XX
@@ 1. Introduction
Contrasting older collections, topics in modern test collections often do not have descriptions or narratives, which is known to affect agreement (Rees and Schultz, 1967), documents are often shorter, and relevance is assessed on a graded scale, which has not been investigated in re-annotation studies. Table 1 contrasts previous studies and the properties of the corresponding test collections with the TREC Deep Learning 2019 (Craswell et al., 2019) (DL’19) test collection we use in our reproducibility work. The changes in modern collections potentially increase assessment variability because an assessor can freely interpret the query intent and context that may be missing from the document to assess its relevance on a fine-grained scale. In this reproducibility study, we investigate if the robustness of classical test collections still holds for modern test collections under these new conditions.
@@
These shifts in test collection design coincide with advances in retrieval methodologies; retrieval pipelines increasingly comprise components where neural language models estimate relevance. These systems have continuously improved over multiple iterations through improved training processes (Pradeep et al., 2022), through distillation from large language models (Pradeep et al., 2023; Schlatt et al., 2024b), or by directly using large language models for ranking (Sun et al., 2023). We hypothesize that iterative model improvements on static benchmarks may overfit to the specific opinions expressed in static relevance judgements, introducing systemic bias, which can be characterized as “bias by research design” in which the influence of previous works can bias new findings implicitly towards a common benchmark (Hovy and Prabhumoye, 2021). The TREC DL’19 track remains a commonly used standard for benchmarking systems, being cited and most likely evaluated hundreds of times (Cited 32 times at ECIR’24 and SIGIR’24 alone). Craswell et al. (2021) alluded to this point, warning against the iterative design processes employed in ad-hoc neural ranking.
@@
Thus, we look to replicate the classic setting for evaluating the robustness of collections proposed by Voorhees (2000) for modern collections, with the additional hypothesis that bias in the system orderings may be introduced through model iteration instead of human disagreement. Since a single expert human assessor’s relevance grades on a given topic are subjective and “equally plausible” as another assessor’s, we can compare the effectiveness of systems using multiple “equally correct” sets of relevance judgments. If a system’s relative effectiveness degrades when evaluated using alternative judgments, it may indicate that models overfit the preferences expressed by the collection’s official judgments.
@@
In this paper, we conduct a re-annotation of TREC DL’19 using two “secondary” annotators for every topic. Based on our new relevance judgments, we first provide a qualitative analysis of this process and validate the findings of previous work in a modern setting. That is, we observe low agreement among assessors but system rankings remain stable. Secondly, treating human annotators as manual TREC runs, we find that one can no longer meaningfully discriminate between the state-of-the-art and human judges on this collection. A reasonable “upper bound” effectiveness is around nDCG@10 of 0.81 (due to the assessment ambiguity). We suggest that reported results that are higher than this bound on DL’19 might already overfit to the relevance labels. Note that we are not aware of any existing system reporting results exceeding this bound. [^1] We do not consider large ensemble systems, given their risks (Amatriain and Basilico, 2017). However, the leading systems today are very close, with RankZephyr (Pradeep et al., 2023) and RankGPT (Sun et al., 2023) reaching between 0.78–0.80 nDCG@10 with a sufficiently strong first stage. We additionally find that closed-sourced large language models and systems distilled from them can fluctuate greatly in relative effectiveness, often degrading system rank on our new relevance judgments.
@@ 2. Related Work
This section outlines evaluation frameworks in retrieval, previous efforts to assess collection reliability, and judgement stability.
@@
The reliability of offline test collections is contested due to the intrinsic subjectivity of relevance under different human annotators (Taube, 1965; Regazzi, 1988), annotation settings (Rees and Schultz, 1967), relevance grading (Cuadra and Katter, 1967) and, more recently, the process of collecting candidate documents via pooling (Zobel, 1998). As these concerns have been raised, one method to validate offline test collections is to re-annotate judgements. Lesk and Salton (1968) and Burgin (1992) empirically validated the collection reliability under annotator disagreement, observing that system order is maintained under different annotators by measuring variance in precision and recall. When web-scale corpora rendered exhaustive annotation impossible, depth pooling became popularized to provide a minimal set of documents that should be annotated to discriminate system effectiveness. Concerns were two-fold: that systems that contributed to the pool would have an advantage over new systems and that the reduced annotations may disrupt the stability of system orderings. Voorhees (2000) found that system order stability could be maintained under the re-annotation of test collection pools from TREC-4 (Harman, 1995) and -6 (Voorhees and Harman, 1997).
@@ 3.1. Reproduction Source
We describe the original annotation process applied by Voorhees (2000). This work was carried out as a component of the curation of the TREC-4 (Harman, 1995) and TREC-6 (Voorhees and Harman, 1997) corpora. As is standard in the curation of pooled TREC corpora, one “primary" annotator was assigned to each topic, subsequently judging each document in the pool. Two additional annotators were assigned to each topic to validate the effect of re-annotation on depth-pooled corpora. A sub-sample of judgements was taken from the pool and re-annotated. In TREC-4, a maximum of 200 relevant documents per topic were taken as a secondary pool for re-annotation. Additionally, 200 non-relevant documents were sampled per topic. A binary judgement was assigned to each query–document pair, relevant or non-relevant. The texts constituting the primary annotation pool, left unjudged in the secondary pool, are included in all system evaluations, maintaining identical pooling biases. The comparison of “secondary" annotations was carried out to assess hypotheses that resurface frequently in IR literature. This work focuses on the subjectivity of relevance and how pooling affects this subjectivity. Thus, experiments were conducted over a depth-pooled corpus under construction at the time. This process produces a subset of judgements where each query–document pair was re-annotated twice.
Where possible, we follow the evaluation process of Voorhees (2000), including analysis of the probability of two systems swapping in effectiveness under different annotators. We define an annotation $a$ as a tuple $\{q,d,r\}$ composed of query $q$, document $d$, and judgement $r\in[0,1,2,3]$. For a given annotator pair and a primary annotator, assigned $n$ query-document pairs, we have three annotation sets $A_{1},A_{2},A_{3}$ where $A_{i}=\{a_{j}\}_{j=0}^{n}$. For $m$ queries, as noted by Voorhees (2000), there are $3^{m}$ possible permutations. Thus, these permutations allow for the greater exploration of judgements that could have been produced from a given annotation process. From the 3 annotation sets, a permutation $A^{*}$ is sampled $s$ times, and systems are evaluated against each permutation. The probability of a swap between systems $i$ and $j$ can be evaluated under a given measure $f$ given judgements $A^{*}$. For systems $S$, we produce a matrix $B\in\mathbb{Z}^{|S|\times|S|}$ where $B[i,j]$ is the number of annotation samples which caused systems $S_{i}$ to have higher effectiveness than $S_{j}$. The swap probability between these systems can then be computed as $\frac{\text{min}(B[i,j],B[j,i])}{s}$. Due to the symmetric treatment of system comparisons, this computation is bounded between 00 and $0.5$.
@@ 3.2. Re-Annotation of a Modern Collection
We treat the original judgements of DL’19 as those of the “primary” annotator. In our study, two annotators from a pool of 8 retrieval academics are assigned to each topic and act as “secondary” annotators aligning with the original study in that each topic is annotated by 2 new annotators; however, the original study does not specify the number of total annotators. The size of the annotator pool used in this work is similar to that of TREC collections (Soboroff, 2024), albeit with a reduced total annotation budget due to cost and following in the sampling setting of Voorhees (2000). We create a secondary pool composed of all documents in the pool with a grade of 1 or more and sample 100 documents per annotator from the non-relevant documents of the primary pool. We reduce the number of non-relevant judgments from the pool motivated by the hypothesis that agreement is far greater over what is non-relevant; for set-based measures, this choice leads to a balance between relevant and non-relevant texts and has minimal effect on nDCG.
Our process for pool creation uses all judged documents at least related to a topic and assigns topics to annotators in a round-robin fashion to balance annotation load while maintaining two annotators per topic. A priority queue of annotator pairs ordered by their current annotation load is produced and assigned the next topic with the greatest number of judgements. The load on each annotator is then balanced by swapping large topics to annotators with minimal load in a second round-robin. We then assign non-relevant texts to each annotator pair with a total mean annotation load of 1125 pairs per annotator.
@@ 3.3. Further Investigation
Judgement Aggregation
Voorhees (2000) apply several aggregation operations including permutation as described in Section 3.1 and aggregate operators. We perform maximum, minimum, and mean pooling over judgments. In the case of mean pooling, we do not perform further quantization; for example, three texts of grades 1, 1.5, and 2 where the second text was annotated 1 and 2 by different annotators would be compared in that order.
@@
To ablate this factor, we fix narratives after the curation of judgements to assess how narrative interpretation can affect agreement and downstream measures of effectiveness. We take three topics of low, median, and high agreement measured via Fleiss $\kappa$ between the primary and secondary annotators; half of the annotators reference one narrative, and half reference the other. We can then further isolate sources of disagreement to query intent and intrinsic ambiguity of a query title both qualitatively and quantitatively.
@@
User models underly many common IR measures such as discounted cumulative gain (Järvelin and Kekäläinen, 2002) and mean reciprocal rank (@$k$(Chapelle et al., 2009). In performing offline evaluations of search systems using these measures, we aim to model effectiveness via user behavior approximated by relevance judgments. In principle, an optimal system would approach a score of 1 for a normalized metric (assuming $k$ relevant documents exist). However, given the subjective and variable nature of relevance under different query intents, such effectiveness, though possible, may not be indicative of a truly effective system. Human annotators outside the original annotation setting may represent strong search systems as they have exhaustively annotated available documents to some cutoff $k$. As such, we measure the performance of each annotator under original relevance judgments as an indicator of how non-pooled modern systems are performing relative to a more realistic oracle. Similarly to idealized discounted cumulative gain, we sort by relevance grade post-aggregation. Though graded relevance lacks granularity in contrast to scalar relevance scores, through in-grade permutations (shuffling), we validate that variations in downstream measures of effectiveness are minimal (SD < 0.0001).
@@ 4.1. Experiment Design
Following the re-annotation process of Voorhees (2000), we observe the impact of re-annotation on the DL’19 test collection, using permutations to assess modern system bias.
@@
The DL’19 track primarily focused on the comparison of neural systems in ad-hoc ranking. For equivalence to the re-ranking task, the top 100 texts of all submitted systems from both re-ranking and retrieval tasks were included in our annotation pool. The top 10 texts from each of the 75 systems (44% Neural LM, 27% Neural, 29% lexical/traditional) were added to our pool, and we use HiCal (Abualsaud et al., 2018) to collect an additional 100 texts per topic.
@@
Voorhees (2000) measure inter-annotator agreement in a pair-wise setting using the overlap of judgements defined as the number of common judgements divided by the total number of judgements. We additionally use Cohen’s $\kappa$ over both 4-grade and binarized relevance and Fleiss’ $\kappa$, allowing for comparisons between 3 annotators (1 primary, 2 secondary) simultaneously. In the TREC-4 and 6 test collections, macro-averaged precision was the primary measure of effectiveness, and therefore, the previous study measured systems by MAP. The TREC DL tracks with four grades of relevance instead primarily measure nDCG@10, and thus, our study is conducted over this metric. Aligned with Voorhees (2000), we measure downstream system order correlation between judgement sets using measure Kendall’s $\tau$; additionally, we measure Spearman’s $\rho$ and RBO.
@@
As we have multiple annotators per query, we can observe pooled judgments across several hypothetical sets; we take all possible combinations of annotators and evaluate both pooled and new models to observe the stability of system order under new judgments. For a given judgment set, we produce a ranking of systems by a target metric (i.e., nDCG@10) and compare this system order to the system ordering under the original judgments. We aggregate rank $\Delta$ for all possible natural combinations over each system to observe how frequently a system degrades over other assessed systems. We measure significant differences in retrieval effectiveness via a paired t-test with Bonferroni correction for different annotator combinations. We measure significant changes in system rank by a Wilcoxon signed test comparing annotator combinations with the system order by official TREC judgements.

### v2

@@ Abstract.
ccs: Information systems Evaluation of retrieval resultscopyright: acmcopyrightjournalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730308isbn: 979-8-4007-1592-1/2025/07
[…]
by
@@ 1. Introduction
Contrasting older collections, topics in modern test collections often do not have descriptions or narratives (Craswell et al., 2019), which is known to affect agreement (Rees and Schultz, 1967; Voorhees and Harman, 1997), documents are often shorter, and relevance is assessed on a graded scale, which has not been investigated in re-annotation studies. Table 1 contrasts previous studies and the properties of the corresponding test collections with the TREC Deep Learning 2019 (Craswell et al., 2019) (DL’19) test collection we use in our reproducibility work. The changes in modern collections potentially increase assessment variability because an assessor can freely interpret the query intent and context that may be missing from the document to assess its relevance on a fine-grained scale. In this reproducibility study, we investigate if the robustness of classical test collections still holds for modern test collections under these new conditions.
@@
These shifts in test collection design coincide with advances in retrieval methodologies; retrieval pipelines increasingly comprise components where neural language models estimate relevance. These systems have continuously improved over multiple iterations through improved training processes (Pradeep et al., 2022), through distillation from large language models (Pradeep et al., 2023; Schlatt et al., 2024b), or by directly using large language models for ranking (Sun et al., 2023). We hypothesize that iterative model improvements on static benchmarks may overfit to the specific opinions expressed in static relevance judgments, introducing systemic bias, which can be characterized as “bias by research design” in which the influence of previous works can bias new findings implicitly towards a common benchmark (Hovy and Prabhumoye, 2021). The TREC DL’19 track remains a commonly used standard for benchmarking systems, being cited and most likely evaluated hundreds of times (Cited 32 times at ECIR’24 and SIGIR’24 alone). Craswell et al. (2021) alluded to this point, warning against the iterative design processes employed in ad-hoc neural ranking.
@@
Thus, we look to replicate the classic setting for evaluating the robustness of collections proposed by Voorhees (2000) for modern collections, with the additional hypothesis that bias in the system orderings may be introduced through model iteration in addition to human disagreement. Since a single expert human assessor’s relevance grades on a given topic are subjective and “equally plausible” as another assessor’s, we can compare the effectiveness of systems using multiple “equally correct” sets of relevance judgments. If a system’s relative effectiveness degrades when evaluated using alternative judgments, it may indicate that models overfit the preferences expressed by the collection’s official judgments.
@@
In this paper, we conduct a re-annotation of TREC DL’19 using two “secondary” annotators for every topic. Based on our new relevance judgments, we first provide a qualitative analysis of this process and validate the findings of previous work in a modern setting. That is, we observe low agreement among assessors but system rankings remain stable. Secondly, treating human annotators as manual TREC runs, we find that one can no longer meaningfully discriminate between the state-of-the-art and human judges on this collection. A reasonable “upper bound” effectiveness is around nDCG@10 of 0.81 (due to ambiguities in relevance assessment). We suggest that reported results that are higher than this bound on DL’19 may already overfit to the relevance labels. Note that we are not aware of any existing system reporting results exceeding this bound. [^1] We do not consider large ensemble systems, given their risks (Amatriain and Basilico, 2017). However, the leading systems today are very close, with RankZephyr (Pradeep et al., 2023) and RankGPT (Sun et al., 2023) reaching between 0.78–0.80 nDCG@10 with a sufficiently strong first stage. We additionally find that closed-sourced large language models and systems distilled from them can fluctuate greatly in relative effectiveness, often degrading system rank on our new relevance judgments.
@@ 2. Related Work
This section outlines evaluation frameworks in retrieval, previous efforts to assess collection reliability, and judgment stability.
@@
The reliability of offline test collections is contested due to the intrinsic ambiguity and biases present in relevance assessment under different human annotators (Taube, 1965; Regazzi, 1988), annotation settings (Rees and Schultz, 1967), relevance grading (Cuadra and Katter, 1967), and, more recently, the process of collecting candidate documents via pooling (Zobel, 1998). As these concerns have been raised, one method to validate offline test collections is to re-annotate judgments. Lesk and Salton (1968) and Burgin (1992) empirically validated the collection reliability under annotator disagreement, observing that system order is maintained under different annotators by measuring variance in precision and recall. When web-scale corpora rendered exhaustive annotation impossible, depth pooling became popularized to provide a minimal set of documents that should be annotated to discriminate system effectiveness. Concerns were two-fold: that systems that contributed to the pool would have an advantage over new systems and that the reduced annotations may disrupt the stability of system orderings. Voorhees (2000) found that system order stability could be maintained under the re-annotation of test collection pools from TREC-4 (Harman, 1995) and -6 (Voorhees and Harman, 1997).
@@ 3.1. Reproduction Source
We describe the original annotation process applied by Voorhees (2000). This work was carried out as a component of the curation of the TREC-4 (Harman, 1995) and TREC-6 (Voorhees and Harman, 1997) corpora. As is standard in the curation of pooled TREC corpora, one “primary" annotator was assigned to each topic, subsequently judging each document in the pool. Two additional annotators were assigned to each topic to validate the effect of re-annotation on depth-pooled corpora. A sub-sample of judgments was taken from the pool and re-annotated. In TREC-4, a maximum of 200 relevant documents per topic were taken as a secondary pool for re-annotation. Additionally, 200 non-relevant documents were sampled per topic. A binary judgment was assigned to each query–document pair, relevant or non-relevant. The texts constituting the primary annotation pool, left unjudged in the secondary pool, are included in all system evaluations, maintaining identical pooling biases. The comparison of “secondary" annotations was carried out to assess hypotheses that resurface frequently in IR literature. This work focuses on the subjectivity of relevance and how pooling affects this subjectivity. Thus, experiments were conducted over a depth-pooled corpus under construction at the time. This process produces a subset of judgments where each query–document pair was re-annotated twice.
Where possible, we follow the evaluation process of Voorhees (2000), including analysis of the probability of two systems swapping in effectiveness under different annotators. We define, in an annotation set $A$, an annotation $a\in A$ as a tuple $\{q,d,r\}$ composed of query $q$, document $d$, and judgment $r\in[0,1,2,3]$. For a given annotator pair and a primary annotator, assigned $n$ query-document pairs, we have three annotation sets $A_{1},A_{2},A_{3}$ where $A_{i}=\{a_{j}\}_{j=0}^{n}$. For $m$ queries, as noted by Voorhees (2000), there are $3^{m}$ possible combinations. Thus, these combinations allow for the greater exploration of judgments that could have been produced from a given annotation process. From the 3 annotation sets, a combination $A^{*}$ is sampled $s$ times, and systems are evaluated against each combination. The probability of a swap between systems $i$ and $j$ can be evaluated under a given measure $f$ given judgments $A^{*}$. For systems $S$, we produce a matrix $B\in\mathbb{Z}^{|S|\times|S|}$ where $B[i,j]$ is the number of annotation samples which caused systems $S_{i}$ to have higher effectiveness than $S_{j}$. The swap probability between these systems can then be computed as $\frac{\text{min}(B[i,j],B[j,i])}{s}$. Due to the symmetric treatment of system comparisons, this computation is bounded between 00 and $0.5$.
@@ 3.2. Re-Annotation of a Modern Collection
We treat the original judgments of DL’19 as those of the “primary” annotator. In our study, two annotators from a pool of 8 retrieval academics [^3] These academics contributed to subsequent analysis and write-up in this work. This decision has minimal effect on our findings; nevertheless, we acknowledge this overlap. are assigned to each topic and act as “secondary” annotators (corresponding to 4 topic groups), aligning with the original study in that each topic is annotated by two new annotators; however, the original study does not specify the number of total annotators. The size of the annotator pool used in this work is similar to that of TREC collections (Soboroff, 2024), albeit with a reduced total annotation budget due to cost and following in the sampling setting of Voorhees (2000). We create a secondary pool composed of all documents in the pool with a grade of 1 or more, and sample 100 documents per annotator from the non-relevant documents of the primary pool. We reduce the number of non-relevant judgments from the pool, motivated by the hypothesis that agreement is far greater over what is non-relevant; for set-based measures, this choice leads to a balance between relevant and non-relevant texts and has minimal effect on nDCG.
Our process for pool creation uses all judged documents at least related to a topic and assigns topics to annotators in a round-robin fashion to balance annotation load while maintaining two annotators per topic. A priority queue of annotator pairs ordered by their current annotation load is produced and assigned the next topic with the greatest number of judgments. The load on each annotator is then balanced by swapping large topics to annotators with minimal load in a second round-robin. We then assign non-relevant texts to each annotator pair with a total mean annotation load of 1125 pairs per annotator.
@@ 3.3. Further Investigation
judgment Aggregation
Voorhees (2000) apply several aggregation operations including combination as described in Section 3.1 and aggregate operators. We perform maximum, minimum, and mean pooling over judgments. In the case of mean pooling, we do not perform further quantization; for example, three texts of grades 1, 1.5, and 2, where the second text was annotated 1 and 2 by different annotators, would be compared in that order.
@@
To ablate this factor, we fix narratives after the curation of judgments to assess how narrative interpretation can affect agreement and downstream measures of effectiveness. We take three topics of low, median, and high agreement measured via Fleiss $\kappa$ between the primary and secondary annotators; half of the annotators reference one narrative, and half reference the other. We can then further isolate sources of disagreement to query intent and intrinsic ambiguity of a query title, both qualitatively and quantitatively.
@@
User models underlie many common IR measures such as discounted cumulative gain (Järvelin and Kekäläinen, 2002) and mean reciprocal rank (@$k$(Chapelle et al., 2009). In performing offline evaluations of search systems using these measures, we aim to model effectiveness via user behavior approximated by relevance judgments. In principle, an optimal system would approach a score of 1 for a normalized metric (assuming $k$ relevant documents exist). However, given the subjective and variable nature of relevance under different query intents, such effectiveness, though possible, may not be indicative of a truly effective system. Human annotators outside the original annotation setting may represent strong search systems as they have exhaustively annotated available documents to some cutoff $k$. As such, we measure the performance of each annotator under original relevance judgments as an indicator of how non-pooled modern systems are performing relative to a more realistic oracle. Similarly to idealized discounted cumulative gain, we sort by relevance grade post-aggregation. Though graded relevance lacks granularity in contrast to scalar relevance scores, through in-grade permutations (shuffling), we validate that variations in downstream measures of effectiveness are minimal (SD < 0.0001).
@@ 4.1. Experiment Design
Following the re-annotation process of Voorhees (2000), we observe the impact of re-annotation on the DL’19 test collection, using combinations to assess modern system bias.
@@
The DL’19 track primarily focused on the comparison of neural systems in ad-hoc ranking. For equivalence to the re-ranking task, the top 100 texts of all submitted systems from both re-ranking and retrieval tasks were included in our annotation pool. The top 10 texts from each of the 75 systems (44% Neural LM, 27% Neural, 29% lexical/traditional) were added to our pool, and HiCal (Abualsaud et al., 2018) is used to collect an additional 100 texts per topic. Our approach does not annotate all texts and thus uses a subset of this pool.
@@
Voorhees (2000) measure inter-annotator agreement in a pair-wise setting using the overlap of judgments defined as the number of common judgments divided by the total number of judgments. We additionally use Cohen’s $\kappa$ over both 4-grade and binarized relevance and Fleiss’ $\kappa$, allowing for comparisons between 3 annotators (1 primary, 2 secondary) simultaneously. In the TREC-4 and 6 test collections, mean-averaged precision (MAP) was the primary measure of effectiveness, and therefore, the previous study measured systems by MAP. The TREC DL tracks with four grades of relevance instead primarily measure nDCG@10, and thus, our study is conducted over this metric. Aligned with Voorhees (2000), we measure downstream system order correlation between judgment sets using measure Kendall’s $\tau$; additionally, we measure Spearman’s $\rho$ and RBO.
@@
As we have multiple annotators per query, we can observe pooled judgments across several hypothetical sets; we take all possible combinations of annotators and evaluate both pooled and new models to observe the stability of system order under new judgments. For a given judgment set, we produce a ranking of systems by a target metric (i.e., nDCG@10) and compare this system order to the system ordering under the original judgments. We aggregate rank $\Delta$ for all possible natural combinations over each system to observe how frequently a system degrades over other assessed systems. We measure significant differences in retrieval effectiveness via a paired t-test with Bonferroni correction for different annotator combinations. We measure significant changes in system rank by a Wilcoxon signed test comparing annotator combinations with the system order by official TREC judgments.